OpenStudy (anonymous):

Can anybody solve this please? Monday night there were 33 kids and 20 adults at the movie theater. That night the box office collected $464 in ticket sales. On tuesday there were 17 kids and 10 adults for a total of $236 in ticket sales. What is the cost for 1 kid's ticket?

OpenStudy (anonymous):

33x + 20y = 464 17x + 10y = 236

OpenStudy (hoblos):

x=cost for 1 kid's ticket y=cost for 1 adult's ticket 33x+20y = 464 17x+10y = 236 ]*2 33x+20y = 464 34x+20y = 472 subtract the 2 equations -x = -8 so x=$8

OpenStudy (anonymous):

1 kid's ticket costs 8 dollars.